1) How would you prepare 500mL of working TBE buffer if the stock solution was purchased as 25X TBE?

2) 2mL of a serum sample was added to 18mL of phosphate buffered saline (PBS) in Tube 1. 10mL of Tube 1 was added to 40mL of PBS in Tube What is the dilution of serum in Tube 2?

3) A student needed to prepare 500mL of 1X TAE buffer to run a QC gel. The stock solution in the lab is 5X TAE. What volumes of stock TAE and water are needed to prepare the 1X working TAE buffer?

4) Four tubes were placed in a rack. To all four tubes, 9mL of PBS buffer was added. 1mL of a protein lysate was added to the first tube and mixed well.   1mL of the first tube is transferred to the second tube and again mixed. The transfer of 1mL from one tube to each subsequent tube continues until 1mL is discarded from the last tube. What is the dilution of protein lysate in the fourth tube?

5) A student labeled 5 tubes 0, 1, 2, 3, 4, 5. Tube 0 contained undiluted protein lysate. The student used the tube 0 lysate to prepared 10-fold serial dilutions in the tubes labeled 1-5. What is the dilution of each of the tubes?

6) A student made an initial 1:5 dilution of protein lysate. Then 2mL of that was added to 8mL of water. Lastly, the student made a 1:20 dilution of the second tube. What is the final dilution of protein lysate.  
Answer #1

Answer 1

As the stock solution is 25x

Working TBE concentration = x

Required TBE volume = 500 mL

Volume of stock =( x*500)/25x = 20 ml

So, 20 ml of stock is to be dissolved in 50 ml of solvent and finally solution must be made to 500 ml.

Answer 2

Dilution factor = final volume/volume of initial solution

Dilution for Tube 1 = (18+2)/2 = 10 times

Dilution for Tube 2 with reference to Tube 1 = (10+40)/10 = 5 times

Dilution of Tube 2 = 5 x 10 = 50 times

Answer 3

Volume of 5X TAE = [(500 ml)*1X]/5X = 100 ml

So, 100 ml of 5X TAE must be added to 400 ml of water.

Answer 4

Dilution of tube 1 = (1+9)/1 = 10 times

In each dilution, 1 ml from previous is diluted to 10 ml.

So, dilution in tube 4 = 10*10*10*10 = 10000 times.

Answer 5

Dilution in Tube 1 = (1+9)/1 = 10

In tube 2, 10*10 = 100

In Tube 3, 100*10 = 1000

In Tube 4, 1000*10 = 10000

In Tube 5, 10000*10 = 100000.

Answer 6

Dilution 1:5 means 1 ml.of sample was added to 4 ml.solvent to make final volume of 5 ml

Dilution 1 = 5/1 = 5 times.

Then 2 ml of tube 1 was added to 8 ml, so, dilution in this step is (2+8)/2 = 5 times

So, till now, dilution is 5*5 = 25 times in tube 2

Then 1:20 dilution means, final volume is 20 and volume of sample from tube 2 is 1 ml

Dilution at this step is 20/1 = 20 times.

Final dilution = 25*20 = 500 times